OpenStudy (anonymous):

Interesting limit: \[ \large \lim \limits_{x\to \infty} \frac{e^{px}}{x^{100}} \]

OpenStudy (lgbasallote):

it's 0! lol

OpenStudy (anonymous):

Nopes.

OpenStudy (anonymous):

Assume p as +ve ;-)

OpenStudy (lgbasallote):

this is inf/inf...let's do LH >:)) e^px(p)/100x^99 still no good... p^2e^px/9900x^98 still no good... this is the circle i have always hated >.<

OpenStudy (anonymous):

Igb I choose this example to show you how to take care of things when L'Hop circles ;)

OpenStudy (jamesj):

You're not in a circle. Just do it 100 times

OpenStudy (lgbasallote):

i cannot find any manipulation here @_@ this is really good

OpenStudy (jamesj):

Apply l'Hopital 100 times and you have \[ \frac{p^100}{100!} e^{px} \]

OpenStudy (jamesj):

\[ \frac{p^{100}}{100!}e^{px} \]

OpenStudy (anonymous):

but there is a smarter way ...

OpenStudy (lgbasallote):

show us oh fool

OpenStudy (anonymous):

lol, think for sometime Igb :)

OpenStudy (jamesj):

Just tell us.

OpenStudy (lgbasallote):

i have thought and i have shared my thoughts.

OpenStudy (lgbasallote):

now it is time you taught us what you have thought

OpenStudy (anonymous):

\[ \large \lim \limits_{x\to \infty} \frac{e^{px}}{x^{100}} = \left( \lim \limits_{x\to \infty} \frac{e^{\frac{px}{100}}}{x} \right)^{100} \] Now apply L'Hop ...

OpenStudy (anonymous):

and you will get your answer \( \infty \)

OpenStudy (nenadmatematika):

just to get involved is p parameter in R?

OpenStudy (jamesj):

I'd like a refund on this one.

OpenStudy (nenadmatematika):

if p>0 than the limit is infinity and if <0 it's zero

OpenStudy (hoblos):

ffm do you know thayler's formula?

OpenStudy (anonymous):

or do you mean taylor's formula?

OpenStudy (hoblos):

lol yeah that one... you can use it

OpenStudy (anonymous):

Yes I know, but I think the way I showed above is easier ...